OpenStudy (anonymous):

Factor. x^2 − 7x + 8 (x + 8)(x – 1) Prime (x – 8)(x − 1) (x + 8)(x + 1)

OpenStudy (anonymous):

I dont understand the question lol

OpenStudy (anonymous):

try multiplying each one out and see if you get the right thing. you won't

OpenStudy (anonymous):

I think it is (x – 8)(x − 1)

OpenStudy (anonymous):

\[ (x – 8)(x − 1) = x^2 -9x +8 \neq x^2 − 7x + 8 \]

OpenStudy (anonymous):

Split the middle term.

OpenStudy (anonymous):

use quardatic formula , if x^2-7x-8 then you are write \

Parth (parthkohli):

\((x + 8)(x – 1) = x^{2} − 7x + 8\) Btw, do you know how to factorize quadratic expressions?

OpenStudy (anonymous):

question is wrong real question is x^2-7x-8

OpenStudy (anonymous):

It's prime.